### v2

# Towards Browser Controls to Protect Cookies from Malicious Extensions

Abstract
Cookies maintain state across related web traffic. As such, cookies are commonly used for authentication by storing a user’s session ID and replacing the need to re-enter credentials in subsequent traffic. These so-called “session cookies” are prime targets for attacks that aim to steal them to gain unauthorized access to user accounts. To mitigate these attacks, the Secure and HttpOnly cookie attributes limit a cookie’s accessibility from malicious networks and websites. However, these controls overlook browser extensions: third-party HTML/JavaScript add-ons with access to privileged browser APIs and the ability to operate across multiple websites. Thus malicious or compromised extensions can provide unrestricted access to a user’s session cookies.
In this work, we first analyze the prevalence of extensions with access to “risky” APIs (those that enable cookie modification and theft) and find that they have hundreds of millions of users. Motivated by this, we propose a mechanism to protect cookies from malicious extensions by introducing two new cookie attributes: BrowserOnly and Monitored. The BrowserOnly attribute prevents extension access to cookies altogether. While effective, not all cookies can be made inaccessible. Thus cookies with the Monitored attribute remain accessible but are tied to a single browser and any changes made to the cookie are logged. As a result, stolen Monitored cookies are unusable outside their original browser and servers can validate the modifications performed. To demonstrate the proposed functionalities, we design and implement CREAM (Cookie Restrictions for Extension Abuse Mitigation) a modified version of the open-source Chromium browser realizing these controls. Our evaluation indicates that CREAM effectively protects cookies from malicious extensions while incurring little run-time overheads.
1 Introduction
Cookies remain a universal state management mechanism on the web. They allow servers to store data on the user’s browser and for that data to be automatically returned in all subsequent requests [1]. This allows servers to link related traffic into a single session and maintain data across the session.
A common and security-critical use of cookies is user authentication, where cookies store an authenticated user session ID and act as a replacement for credentials [2, 3]. Naturally, this makes cookies a valuable target for attacks such as session hijacking, session fixation, and Cross-Site Request Forgery (CSRF). Session hijacking [4, 5, 2] attempts to gain access to the user’s account by stealing the value of their session cookie. These attacks are not limited to session cookies alone as non-session cookies can still leak users’ private information [6]. Similarly, session fixation [3, 5, 7] injects a malicious session cookie with a pre-known value into the user’s browser. Once the user authenticates with the website, an adversary can use the pre-known ID to access the user’s account. Finally, CSRF tricks the browser into issuing a cross-site request to perform some action as the user without their knowledge [3, 8, 5].
Several cookie features have been introduced to combat these attacks [9, 1]. The Secure attribute restricts cookies to HTTPS traffic. Thus Secure cookies are always encrypted on the network. Similarly, the HttpOnly attribute prevents JavaScript (JS) from accessing cookies. Together these attributes help mitigate session hijacking by preventing cookie theft from malicious JS and over the network. The SameSite attribute restricts how cookies are attached to cross-site requests, reducing the threat of CSRF. Cookies also define two cookie prefixes: __Secure- and __Host-. These prefixes enforce that certain criteria be true when creating the cookie. As such, cookie prefixes help reduce the threat of session fixation by asserting how a cookie was created.
These controls reduce the threat posed by malicious websites and networks, however, browsers also support third-party HTM/JS applications called extensions [10, 11]. Extensions add functionality to browsers, run across many websites, and can bypass several browser controls by default, including cookie controls [12, 13]. Despite being JS themselves, extensions are not restricted by HttpOnly. Similarly, extensions can access network traffic as it is processed in the browser. Due to this, they can access cookies within network headers before TLS encryption, bypassing the Secure attribute. Further, extensions can add cookies to traffic, ignoring the SameSite attribute. Extensions can also create cookies with known prefixes and edit cookies/remove their security attributes. As such, extensions still pose a large threat to cookies.
These permissive extension controls stem from the current browser trust model. Existing browser controls often assume two possible attackers, a web attacker [14, 15, 16, 17, 18, 19, 20, 21] and a network attacker [3, 15, 16, 18, 19]. In both models, the attacker controls a malicious website, whereas a network attacker can also inspect and tamper with HTTP traffic. In both models, the attacker is an external party. Thus, everything inside the browser is considered trusted, including extensions. This is also reflected by the extension architecture itself assuming “benign but buggy" extensions [19].
Instead of architectural controls, browser vendors try to prevent malicious extensions from becoming available in the first place. Extensions are reviewed to adhere to a browser’s extension policies [22, 23] before being listed publicly. However, malicious extensions can pass this review [24, 25, 26], requiring them to be pulled from browsers after the fact [27, 28]. Extension permissions must also be granted at installation [29]. Thus users can prevent malicious extensions by denying them permission, however, users are often unaware of the risks associated with these permissions. [30]. Regardless of either check, extensions can also "become" malicious. Extensions update automatically in the background [31]. As such, extensions can silently receive a malicious update while maintaining their permissions [26]. Browsers also allow the manual installation of extensions for development [32]. This can be leveraged to bypass all checks by tricking the user into manually installing a malicious extension.
Prior work also demonstrates that vulnerable extensions can leak their permissions to a malicious page [33, 18, 34]. Thus, compromised extensions can also be used to bypass existing cookie controls. To demonstrate this risk, we use the extension APIs to hijack test accounts on several websites (see Section 3). We also scrape 156,783 extensions from both Chrome’s and Firefox’s extension repositories to assess the prevalence of these APIs in current extensions.
In this work, we aim to reevaluate the trust model employed by current browsers to mitigate risks posed by extensions. Prior work has sought to protect sessions from extensions by preventing their access to sensitive information on web pages [35, 36, 37]. While these controls protect sessions from direct access by extensions, malicious extensions maintain access to session cookies. Thus adversaries can still gain unauthorized access to the session through these cookies.
We propose two new cookie attributes to protect cookies from malicious extensions: BrowserOnly and Monitored. Unlike previous controls, we assume extensions are untrusted and potentially malicious. The BrowserOnly attribute prevents extensions from accessing cookies outright. However, some cookies must remain accessible for websites and extensions to function properly. Thus the Monitored attribute allows extensions to access cookies, ties cookies to a single browser instance, and maintains a changelog for each cookie. This changelog and other server-defined data are authenticated and sent to the server alongside each cookie. While Monitored does not prevent cookie theft, it enables server-side detection of stolen and illegally modified cookies. Note that these features do not imply privacy loss (see Section 8).
The Monitored attribute uses symmetric keys to link cookies to a single browser and authenticate their changelog-s. Specifically, the browser generates a unique key for each site it visits and attaches the key to all outbound HTTPS requests to its respective site. The server then uses the received key to identify the browser by saving the key in the cookie as part of the Monitored attribute. When the server receives the authenticated attribute in later requests, it can verify it using the saved key. This verification ensures that the cookie came from the correct browser and its changelog has not been tampered with. While these keys allow individual sites to identify the user’s browser, they cannot be used to track the user across different sites as the browser uses a unique key for each site.
Support for the BrowserOnly and Monitored attributes is obtained by modifying the browser architecture to revoke an extension’s ability to access/modify these attributes before their usage by the underlying browser core. As a proof of concept, we implement these features as an open-source modified version of the Chromium browser [38] and evaluate our prototype concerning storage and run-time impacts. In summary, our intended contributions are as follows:
  • $\bullet$
    We study Chrome’s and Firefox’s extension repositories to assess the prevalence of extensions that, if exploited, could threaten session cookies.
  • $\bullet$
    We propose a new browser trust model regarding extensions, in which we introduce two new cookie attributes to protect cookies from potentially malicious browser extensions. The BrowserOnly attribute prevents cookies from being accessed by extensions outright. The Monitored attribute leaves cookies accessible while allowing servers to detect cookie theft and tampering (allowing servers to refuse service to stolen cookies). To support this functionality, we design a mechanism to tie browser instances to cookies without linking users across different domains.
  • $\bullet$
    To realize this concept we implement and open-source [39] CREAM (Cookie Restrictions for Extension Abuse Mitigation) as a modified version of Chromium and evaluate its functionality and associated overheads.
2 Background
2.1 Browser Architecture
Modern browsers implement a multi-process design [40, 41, 42] (depicted in Figure 1), leveraging OS-based inter-process isolation to provide strong separation of privilege within the browser. While different browsers have different implementations and naming conventions, all mainstream browsers follow a general model. Namely, they implement two main process types: browser and renderer processes [43, 44, 14].
Renderer processes parse untrusted, potentially malicious, web content (i.e. HTML, CSS, JS) and build the final page displayed to the user [43, 17, 45]. As such, they are unprivileged. Renderers also execute in a sandbox to further restrict their access to the underlying host. This forces all host-related requests to go through the browser process.
The browser process is the main privileged process in the browser [43, 14]. While there can be arbitrarily many renderer processes, each browser instance only has one browser process. The browser process interacts directly with the host, manages all other processes in the browser, and facilitates inter-process communication (IPC) between them [44, 33].
Browsers are also service-oriented [46, 47]. Many renderer and browser process features have been moved to independent services that are often executed within separate processes. For example, browsers have a Network Service that implements the browser’s network stack (all low-level networking code) and handles all network transactions [48, 42, 49]. As these feature-specific processes have no standard name, we use Chrome’s nomenclature of “utility processes” [43, 44] to refer to them. Utility processes can be privileged, unprivileged, sandboxed, or unsandboxed depending on their functionality.
In the remainder of this paper, we often reference “the browser”. When referring to “the browser” we are discussing the collection of features commonly referred to as the browser. This includes both the privileged browser process as well as utility processes, such as the Network Service. We refer to the browser process alone as the “browser core” instead.
Figure 1: The Browser Architecture and Cookie Interfaces
2.2 Extension Architecture
All extensions have a manifest.json file that describes them [11, 19]. Currently, there are two manifest versions (V2 and V3), with slight differences in how attributes are defined and which extension APIs are available. This file lists the APIs and scripts an extension uses, websites it runs on, how it is loaded, resources it can access, and more. The browser requires the manifest to be present when an extension is installed and restricts the extension’s behavior to what is stated in its manifest. This prevents extensions from gaining more privileges after installation, except for optional permissions which can be requested at any time, but must still be declared in the manifest file at installation.
An extension’s functionality is divided into two components: the extension core and content scripts [19, 18]. The extension core is highly privileged with access to the extension APIs and the ability to connect to remote servers. To protect these privileges from untrusted websites, the extension core executes in its own process. Further, as different extensions are effectively independent programs, different extensions’ cores cannot share the same process [50]. However, despite being highly privileged, these processes are also run within a sandbox; again forcing all interactions with the underlying system through the browser [19].
This isolation prevents the extension core from interacting directly with any website. Instead, content scripts are injected into the page and have direct access to its content [19, 18]. These scripts execute within the renderer process of the website and are thus at an increased risk of attack. Due to this, content scripts are isolated from other scripts on the page [19, 33]. Similarly, content scripts are unprivileged and cannot access the extension APIs. Instead, content scripts can only send messages to the extension core using the message-passing API. This aims to prevent a compromised renderer from gaining the powers of any extensions embedded within them.
2.3 Cookies & Cookie Management
At their simplest, cookies are key-value pairs associated with the server that creates them. Cookies can also set multiple optional attributes that further define how they are handled. Which server a cookie belongs to and what requests it should be attached to are determined by the Domain and Path attributes [1]. These indicate what websites and which portion of those sites the cookie can be sent to respectively. If not specified by the server, the browser uses the hostname and path of the website that created the cookie by default. Cookies can also be partitioned. Partitioning limits a cookie’s accessibility to its embedding context such that cookies set by the same site but embedded on different pages are not accessible to each other [51]. This is often used to prevent third-party cookies from tracking users across websites. As discussed in Section 1, cookies also have three security attributes: HttpOnly, Secure, and SameSite. They also support two prefixes that enforce certain criteria when creating a cookie [9].
2.3.1 Cookie Manager
Cookies can be set in many ways. Most commonly, they are defined in the Set-Cookie HTTP header [52] and created by the Network Service [53]. Since cookies provide state for network requests, they are stored within and are directly interactable from the Network Service. Cookies can also be created from a webpage through ‘document.cookie’ [54] or the Cookie Store API [55], by the browser core, and by extensions with the Cookies extension API permission [56]. However, none of these can interact with cookies directly. Instead, the Network Service exposes APIs for accessing its internal cookie store [53]. The arrows in Figure 1 depict how different browser components access cookies.
These interface APIs are implemented by the Cookie Manager and the Restricted Cookie Manager [53]. The Cookie Manager allows unrestricted querying of the cookie store and is used by the browser core. Since the Cookies API is a browser-level API, extensions also access cookies through the Cookie Manager. The Restricted Cookie Manager, as the name implies, is more limited. The restricted manager can only access the cookies of a single specified origin. For this reason, the restricted manager is used by renderer processes to enforce stricter limitations on websites.
2.3.2 Cookie Monster
All cookies are maintained by the browser’s cookie store. Chromium’s concrete implementation of the cookie store is called the Cookie Monster [57]. There are two ways to fetch cookies from the Cookie Monster [58]. Either the Cookie Monster can return all unexpired cookies or a subset of cookies that match some provided criteria. We refer to these methods as GetAll and GetList in the rest of this paper. When using GetAll, the Cookie Monster returns all cookies within its store, regardless of their attributes. However, most fetches use GetList which takes a domain and a set of options, fetches all the provided domain’s cookies, and filters them based on the provided options [53]. These options include checks such as whether to allow HttpOnly or Secure cookies in the results.
To set a cookie (Set operation), the caller (browser core, renderer, extension, etc.) creates a new cookie and sends it to the Cookie Monster along with its domain and other cookie options [58]. Upon receipt of the cookie, the Cookie Monster first checks that the cookie is valid. Similar to GetList, this validation uses the supplied options to ensure the cookie is properly configured (e.g. verifying that an HttpOnly cookie is not being set by JS). If valid, the Cookie Monster then checks for any existing matching cookie. If a match is found, it is deleted and replaced with the new cookie. Otherwise, the cookie is added to the internal map of cookies as a new entry.
3 On the Extent of Extension Threats
3.1 Manipulating Cookies via Extension APIs
To illustrate the practical threats extensions pose to cookies, we implemented three test extensions targeting different extension APIs and tested them in Chrome. Each test extension was built in both manifest versions (V2 and V3), when applicable. We specifically target the Cookies, WebRequest, and DeclarativeNetRequest APIs, as they allow interaction with cookies. The Cookies API allows extensions to get, set, remove, and be notified of changes to cookies within the browser [56]. The cookies accessible to an extension are limited to those of websites it has permission to run on. Nonetheless, an extension can have permission to run on all websites (which is not uncommon). The WebRequest API allows extensions to view network traffic mid-processing [59]. As such, it allows extensions to access HTTP headers, including cookie headers. The WebRequest API can also be extended with the WebRequestBlocking permission to allow extensions to pause the execution of network events and edit request headers. Finally, the DeclarativeNetRequest API also allows extensions to modify network requests [60, 61]. However, unlike WebRequest, DeclarativeNetRequest cannot view the contents of the request. Each test extension targeted a single API as follows:
  • $\bullet$
    Case 1 - Cookies API Extension: This extension implements a cookie management interface allowing the user to view, edit, and delete browser cookies. When fetching cookies the extension also forwards them to a malicious server. Using this we stole the cookies of many websites including Google, Amazon, and Facebook. Adding the stolen cookies to a new Firefox instance allowed us to access the corresponding test accounts without their credentials. We note that these accesses require a combination of cookies rather than a single cookie, but as extensions can easily access all of a website’s cookies this is not an obstacle.
  • $\bullet$
    Case 2 - WebRequest API Extension: This extension allows the user to intercept, view, and modify traffic for specified websites. When viewing the request, the request headers are also forwarded to a malicious server. From the headers, we can steal all the cookies sent to and set by the website. This extension can also modify request headers and thus edit the cookies sent to/received from the server. It is important to reiterate that modifying headers requires the WebRequestBlocking permission [59] which is only available for V2 extensions in Chrome. While V3 extensions cannot edit requests they can still steal cookies.
  • $\bullet$
    Case 3 - DeclarativeNetRequest API Extension: This extension allows the user to create rules that modify network traffic without viewing its content. As such, it allows us to edit the cookies in traffic, however, as the extension cannot inspect the request headers, it cannot steal cookies. The Declarative API only exists for V3 extensions.
Table 1: Breakdown of extensions versioning, host permissions, and access to cookies
Browser Manifest-Version Host Permissions APIs
V2 V3 $<$all_urls$>$ https://*/* http://*/* Cookies WebRequest (WR) WebRequestBlocking (WRB) DeclarativeNetRequest (DNR)
Firefox 34,929 2,702 13,180 221 57 2,154 5,219 3,626 66
Chrome 59,020 60,132 40,363 1,452 268 5,728 6,620 3,352 2,048
Table 2: Total users affected by each API
Browser Cookies WR WRB DNR
Firefox 13,909,278 47,092,558 43,306,429 70,374
Chrome 408,937,955 947,610,948 489,938,433 353,664,323
Table 3: Extensions with "$<$all_urls$>$" and cookie access
Browser Cookies WR WRB DNR DNR + WR
Firefox 1,409 3,440 2,415 44 7
Chrome 3,274 4,278 2,032 1,391 299
Most modern browsers are implemented on top of a few open-source browsers. Chrome, Microsoft Edge, Opera, and Brave are all built on top of the Chromium browser [38]. As such, our extensions worked on all four browsers without modification. We also tested the extensions in the Firefox-based browsers: Firefox and Waterfox. The V2 extensions (Cookies and WebRequest) were directly transferable as Firefox also implements the WebExtension APIs [62] to support cross-browser extensions. When using V3, one change was required. Despite requiring that host permissions are defined in an extension’s manifest, Firefox treats them as optional permissions for V3 extensions [63]. Thus, after installation, the extension must prompt the user to re-accept its host permissions. Once granted, V3 extensions behaved as they did in Chromium. We believe that prompting the user does not pose a problem to the adversary, as they could (and likely would) imitate a legitimate service. Another caveat is that Firefox still supports the WebRequestBlocking permission for V3 extensions. Testing in Waterfox obtained the same results.
3.2 Extension API Permissions’ Landscape
The above examples highlight how certain extension APIs can be leveraged to steal and modify cookies (including session cookies). To determine the prevalence of these permissions we crawled Chrome’s and Firefox’s extension repositories: the Chrome Web Store [64] and addons.mozilla.org (AMO) [65]. As of March 2024, we identified 122,951 Chrome and 37,783 Firefox extensions. However, of the extensions identified, we could only pull the manifests of 119,152 Chrome and 37,631 Firefox extensions for analysis. Appendix A discusses the scrape and identified extensions in greater detail. We also identified several Chrome Apps and Android extensions which we discuss in Appendix B. All extensions discovered and manifests analyzed are available in our public repository [39].
We parsed each extension’s manifest for three attributes. First, we obtained the manifest version from the manifest_version tag. Next, we checked the permissions entry to find if the extension used any of the three APIs discussed in Section 3.1. Finally, we searched the manifest for the strings “$<$all_urls$>$”, “*://*/*”, “https://*/*”, and “http://*/*”. These strings allow an extension to access every page (first two strings), all HTTPS pages, and all HTTP pages in the browser, respectively [66]. As these host permissions can appear in various places, their presence indicates that at least a portion of the extension has this access. Similarly, not all extensions with “$<$all_urls$>$” behavior use this tag. Due to this, we treat “$<$all_urls$>$”, “*://*/*”, and the combination of “https://*/*” and “http://*/*” as equivalent. Table 1 summarizes our results. We also calculate the number of users affected by each API as reported by each extension store (depicted in Table 2).
We found that 59.9% of extensions are written using manifest V2. However, V2 extensions are much more prevalent in Firefox accounting for 92.8% of all Firefox extensions. Chrome extensions are more evenly split with 49.5% still using the V2 format. This disparity is likely due to Chrome’s plan to deprecate the V2 format in 2025 [67] whereas Firefox will continue to support both versions [68].
For host permissions, we found that 35.2% of the extensions analyzed define broad host permission. Unlike manifest versioning, this ratio is consistent across Firefox (35.6%) and Chrome (35.3%). Nearly all of these extensions have the “$<$all_urls$>$” behavior. This accounts for 34.2%, 35%, and 33.9% of all, Firefox, and Chrome extensions, respectively. 0.2% of extensions declared permission for all HTTP sites only and 1% declared permission for all HTTPS sites alone.
7,882 extensions (5% of extensions analyzed) used the Cookies API, reaching a combined user base of 422,847,233 users. Note that this combined user base represents an upper bound, as a single user may have multiple extensions and thus be counted multiple times. Similarly, it is unclear whether extension repositories measure users by installation or actual user accounts. More extensions used the WebRequest API. Specifically, 11,839 (7.6%) were able to monitor network traffic and 6,978 extensions (4.5%) also declared the WebRequestBlocking permission and could edit HTTP headers. Again while limited extensions declare these permissions, they have nearly one billion users (994,703,506) with 533,244,862 users simultaneously affected by the WebRequestBlocking permission. The DeclarativeNetRequest API is the least used, only appearing in 2,114 extensions (1.3%). Similarly, it is almost non-existent in Firefox appearing in only 66 extensions (0.17% of all Firefox extensions). This low adoption is likely due to the ongoing transition from V2 to V3 formatting (at least in Chrome), as the DeclarativeNetRequest API is only available to V3 extensions. In that regard, it appeared in 3.4% of V3 extensions found and has a combined 353,734,697 users.
Each API is limited by the extension’s host permissions. Hence, we also examined how many of the above extensions also have “$<$all_urls$>$” host permissions. Table 3 summarizes these findings. 54.8% of extensions with Cookies API permissions have access to all cookies in the browser. Similarly, 65.2% of WebRequest extensions and 63.7% of WebRequestBlocking extensions can read and modify any requests in the browser. 67.9% of manifests with DeclarativeNetRequest permissions can also modify any network requests. We also found that 21.3% of these extensions (14.5% all declarative extensions) also contain WebRequest permissions, allowing them to view any request (despite contradicting the privacy-preserving philosophy of the declarative API).
From our analysis, we find that widespread extensions with a combined user base in the order of hundreds of millions can access every cookie in the user’s browser. When compromised (or malicious), these extensions pose a significant threat to cookies and session security.
4 CREAM Overview
Motivated by the results presented in Section 3, we advocate for browser controls to allow selective prevention or detection (when prevention is not possible due to application requirements) of cookie theft and illegal cookie modifications made by extensions. To realize these goals we propose CREAM (Cookie Restrictions for Extension Abuse Mitigation) as a modification to the trust model within browsers. This Section focuses on CREAM’s concepts and protocol. We discuss browser support for CREAM in Section 5.
4.1 Threat Model
We extend the active network attacker model described in prior work [15, 16, 19]. The model assumes an Adversary ($\sf{\mathcal{A}dv}$) can intercept, modify, and inject traffic on the network. It also assumes $\sf{\mathcal{A}dv}$ controls a malicious website. In addition, we assume that $\sf{\mathcal{A}dv}$ controls a malicious extension in the user’s browser with arbitrary API permissions and can run on all pages in the browser (e.g. has the “$<$all_urls$>$" host permission). As such, $\sf{\mathcal{A}dv}$ can also inject content scripts (malicious JS) into all pages. $\sf{\mathcal{A}dv}$’s goal is to modify or steal a user’s cookies through attacks such as cookie hijacking and cookie fixation and gain access to the user’s session.
We assume that the browser core and the Network Service utility process are trusted and vulnerability-free, as CREAM functionality is implemented as part of these modules. Similarly, we trust the underlying host’s OS which is relied upon to implement all browser isolation mechanisms (irrespective of CREAM). Attacks that compromise/modify privileged software (e.g., the OS) or the browser software implementation are considered orthogonal and out of scope in this paper.
4.2 Rationale
CREAM introduces two new cookie attributes: BrowserOnly and Monitored. BrowserOnly is a 1-byte flag indicating whether a cookie should be visible to extensions. If set, the cookie is only accessible to the browser. Therefore neither JS nor extensions can access these cookies. The BrowserOnly attribute only protects cookies within the browser. To prevent BrowserOnly cookies from being stolen on the network they should also be marked Secure, which is standard.
For cases where cookies must be accessible by websites and extensions (where BrowserOnly can not be used), we define the Monitored attribute. The Monitored attribute holds a server-defined message that when received is stored by and only accessible to the browser. The browser also logs all changes made to Monitored cookies. Later, when sending a Monitored cookie to the server, the browser attaches an authenticated report to the request that contains this server-defined message, the cookie’s current settings, and any logged changes. With this, the server can validate the changes made to the cookie and detect if it is stolen.
Since Monitored cookies must be accessible, we cannot prevent them from being stolen. Instead, we make stolen cookies unusable (e.g., for authentication). Figure 2 depicts the protocol to create (part (A)) and use (part (B)) Monitored cookies. The browser generates a symmetric key for each site it visits (step A.1 in Figure 2) and attaches it to all outbound HTTPS requests for that site (step A.2). Note that in this context, a site refers to the protocol and registry controlled domain name of a resource [17, 69] (e.g., https://example.com). The generated key allows a site to identify the browser in subsequent communication. However, by creating a new key per site these keys cannot be used to track users across different sites. When creating a Monitored cookie, the server builds the Monitored attribute’s message. This message must include the cookie’s name, value, and the received browser key. The cookie’s name-value pair ties the attribute to this specific cookie while the browser key ties the whole cookie to a single browser. Along with these values, the server can include any other data necessary for later cookie validation. The server then uses a secret to symmetrically encrypt the message and uses the resulting ciphertext as the Monitored attribute’s value (step A.3). This authenticated encryption ensures the attribute was generated by the server, is only readable to the server, and cannot be altered.
Server Browser
(A) CREATING A Monitored COOKIE
1). $K_{browser}\leftarrow KeyGen()\qquad$
2). Attach $K_{browser}$ to outbound request
[…]
3). $msg=Enc((Name,Value,K_{browser},...),K_{server})$
      ($K_{server}$ is a secret known only by the server)
[…]
4). Store the received cookie
(B) SERVER RECEIVING A Monitored COOKIE
1). Fetch cookies for the domain
2). If cookie $C_{i}$ has the Monitored attribute:
       $Report_{C_{i}}=\{log_{C_{i}},Settings_{C_{i}},msg_{C_{i}},timestamp\}$
       H $=HMAC(Report_{C_{i}},K_{browser})$
3). Attach $Report_{C_{i}}$ and H to the outbound request
[…]
4). $Name_{exp},Value_{exp},K_{browser},...\leftarrow Dec(msg_{C_{i}},K_{server})$
5). Check: $Name_{C_{i}}==Name_{exp}~{}\&\&~{}Value_{C_{i}}==Value_{exp}$
6). $Verify(Report_{C_{i}},H,K_{browser})$
7). Use $Report_{C_{i}}$ to validate $C_{i}$ (per server policy)
Figure 2: Protocols for creation and usage of Monitored cookies
When sending the Monitored cookie back to the server, the browser builds its corresponding Monitored report, including an HMAC produced with the generated key for that site, and attaches it to the request (steps B.1-B.3). Upon receipt, the server decrypts the Monitored message with its secret key (B.4), validates that it is the correct message for the received cookie (B.5), and verifies the HMAC using the key stored in the Monitored attribute (B.6). If the cookie was stolen by a malicious extension, the HMAC verification would fail as $\sf{\mathcal{A}dv}$ would not know the browser’s key (as long as the browser is itself trusted – recall Section 4.1). Similarly, attempts to change the server-defined Monitored message would cause the decryption/validation to fail. As a result, Monitored cookies can only be used by the legitimate browser.
As detection relies on the symmetric key sent to the server in the initial request, the key can only be sent over secure (HTTPS) channels. This ensures the key is always encrypted and inaccessible to $\sf{\mathcal{A}dv}$ over the network. Otherwise, if the key was not encrypted $\sf{\mathcal{A}dv}$ could steal it to forge authenticated Monitored reports and use stolen Monitored cookies. However, once established, Monitored cookies can be sent over secure and insecure channels as the key in the Monitored attribute is always encrypted by the server’s key.
By design, Monitored cookies are editable. As such, the browser logs all changes made to Monitored cookies. This changelog allows the server to validate these changes and take action accordingly (depending on the server’s policy). While the changelog stores all changes made to the cookie, it only exists while the cookie exists. Therefore, if $\sf{\mathcal{A}dv}$ (in the form of an extension) deleted the cookie and made a new cookie shadowing the original [5] except for $\sf{\mathcal{A}dv}$ intended changes, these changes would not be logged as no changelog exists at this point. Similarly, only edits made through the Cookie Monster are logged thus changes made directly to cookie headers are not. Therefore, when sending the Monitored report to the server, CREAM includes the current value of the cookie’s attributes (i.e. its current settings) as well as the changelog. By comparing these current settings to some expected value, the server can detect changes that do not appear in the changelog. The server-authenticated Monitored attribute can store the expected settings if needed. Finally, CREAM also adds a timestamp in the report to prevent replay attacks. As stated above, the entire report is authenticated and sent to the server alongside the Monitored cookie.
Leveraging these controls requires servers to update their behavior, adding the attributes to cookie definitions and, in the case of Monitored, implementing the discussed protocol accordingly. However, these attributes do not alter existing cookie attributes and behavior. Thus only servers that wish to use the BrowserOnly or Monitored attributes must alter their behavior, preserving backward compatibility.
5 CREAM Architectural Design
As discussed above, the BrowserOnly and Monitored attributes require architectural support from the browser to restrict/monitor extension actions. Aside from creating the new cookie attributes, this also requires support from the Cookie Monster and Network Service. As each attribute functions independently, we discuss their required support separately.
(A) Outbound Request
(B) Inbound Response
Figure 3: The flow of requests and responses are through the Network Service
5.1 BrowserOnly Attribute
As discussed in Section 4, BrowserOnly is a 1-byte flag that prevents a cookie from being accessed by JS and extensions. Thus the Cookie Monster must be updated to filter on this attribute. Recall from Section 2.3.2 that there are two ways to fetch cookies from the Cookie Monster, GetAll and GetList, and both GetList and Set filter cookies based on a supplied set of options. As such, we create a new option to exclude BrowserOnly cookies from GetList and Set calls. This option prevents BrowserOnly cookies from being returned or overwritten in most cases. However, as GetAll never filters its results, BrowserOnly cookies are still returned. To address this, we create a new endpoint (GetAllExt) that returns all non-BrowserOnly cookies in the Cookie Monster. These changes are then propagated to the Cookie Manager API and its respective implementations. Finally, the Cookies extension API is updated to use GetAllExt instead of GetAll and the option to exclude BrowserOnly is added to the necessary accesses. This blocks all extension and JS accesses to BrowserOnly cookies through the Cookie Monster. Note that cookie deletion in response to storage restrictions (“purges”) can be conditioned on cookies’ priorities and security attributes [58]. Therefore, we also configure the Cookie Monster to not evict BrowserOnly cookies in these scenarios, thus mitigating eviction attacks [5, 70].
While the above changes make BrowserOnly cookies inaccessible to the Cookies API (and JS), they are still exposed to the WebRequest and DeclarativeNetRequest APIs. Protecting BrowserOnly cookies from these APIs requires updating how cookies are handled in network traffic. Figure 3(A) illustrates how the Network Service builds network requests. When navigating to a website, the Network Service creates a URLRequestHTTPJob to handle the request (step A.1 in Figure 3(A)). When building the request, the HTTPJob fetches the destination’s cookies from the Cookie Monster and adds them to the Cookie request header (A.2 & A.3). When all the request headers have been created, they are passed to extensions with the above APIs for further processing (A.4). Once all extensions have finished, the headers are returned to the HTTPJob (A.5) and the request is sent to the server (A.6). To hide BrowserOnly cookies from extensions, we withhold them from the Cookie header until after the extensions have executed (A.5). Once the request headers are returned, we add the BrowserOnly cookies to the front of the Cookie header. This ensures the BrowserOnly cookies are parsed first and prevents cookie shadowing attacks [5].
Similarly, Figure 3(B) displays how the Network Service parses the server’s response. When the response is received (step B.1 in Figure 3(B)), the URLRequestHTTPJob forwards the response headers to extensions (B.2). Once the extensions finish, the headers are returned to the HTTPJob (B.3). At this point, any new cookies are added to the Cookie Monster (B.4) and the response is sent to the renderer for display (B.5). To hide BrowserOnly cookies in responses from extensions, we strip all Set-Cookie headers with the BrowserOnly attribute from the response on arrival (B.1). Then when the headers are returned from the extensions (B.3) we reattach the BrowserOnly headers to the response.
For simplicity, this explanation omits some details of how the Network Service handles transactions and interacts with extensions. Requests pass through several classes between the renderer and the network. We focus on the URLRequestHTTPJob as all BrowserOnly (and Monitored) functionality was added to this class. Similarly, when building a network request, its headers are sent to extensions up to three times [59]. Protecting BrowserOnly cookies requires adding the cookies to the request only after all extension passes have finished. The same applies to parsing network requests.
Remark: While BrowserOnly functionality could in principle be incorporated into existing security-relevant attributes (e.g., HttpOnly) merging these functions would lead to changes to the standard cookie usage model. CREAM instead creates a new attribute that can be used independently, preserving existing usage models and backward compatibility.
5.2 Monitored Attribute
The Monitored attribute instructs the browser to maintain a server-defined message and a changelog for the cookie. These values are only accessible to the Network Service. As such, Monitored only requires the Cookie Monster’s Set behavior be updated to log changes to Monitored cookies. As discussed in Section 2.3.2, Set takes a new cookie, checks if a matching cookie already exists, and replaces it if found. Hence when setting a cookie, if a preexisting match is found and it is a Monitored cookie, we check whether the new cookie has a Monitored attribute. If it does not, the matching cookie’s Monitored message and changelog are copied to the new cookie. Then the two cookies are compared and any differences are added to the newly copied changelog. If the new cookie has a Monitored attribute, the new attribute’s message is kept and the changelog is cleared. Similar to BrowserOnly, extensions (and JS) cannot create Monitored cookies. Thus, a cookie with a Monitored attribute must be set by the server and all existing changes must have been seen by the server. This behavior allows the server to update the Monitored attribute as needed and minimize the size of the changelog by removing previously received/validated entries. To avoid resource exhaustion attacks [71], a maximum number of changelog entries can be set. The cookie can simply be flagged as invalid upon reaching the limit.
As discussed in Section 4, along with the Monitored attribute, the browser generates a symmetric key for each site it visits and attaches the key to all outbound HTTPS requests to that site. The browser must maintain a map of keys and their associated sites that is available to URLRequestHTTPJobs. These keys must be created by and maintained within the Network Service. When sending a request, the URLRequestHTTPJob simply fetches the correct key from this map and attaches the key to the request using a custom header. To avoid exposure/tampering by extensions, the key is added after extensions have processed the request headers.
When adding Monitored cookies to network requests (or whenever fetched in general), the changelog and Monitored message are never included in the cookie. As such, Monitored cookies can be seen by extensions. However, when building the Cookie header, we log which cookies are Monitored cookies. Then, after the extensions have processed the headers, we generate the Monitored report for the Monitored cookies. Specifically, for each Monitored cookie, we retrieve the cookie’s changelog, current settings, and Monitored message from the Cookie Monster and add them to the report. Once all the cookies have been added to the report, the report is timestamped and authenticated with the generated key for the site. The report is authenticated as a whole rather than per individual cookie’s sub-report because all cookies of a single site use the same key. Nonetheless, each cookie contains the key within its authenticated/encrypted Monitored message to ensure that the server always receives (and can verify) the key regardless of which set of cookies is present.
Handling Monitored cookies on incoming responses requires no additional changes. Unlike BrowserOnly cookies, Monitored cookies can be seen by extensions. Similarly, even if extensions see the Monitored message, they cannot change it or use it outside of the user’s browser. At best, an extension can remove the attribute and disable the changelog. However, this would only lead the server to reject the cookie as it would not contain the expected report.
6 Security Analysis
We argue that the BrowserOnly and Monitored attributes protect cookies against malicious extensions (and JS), including those with highly privileged API permissions (analyzed in Section 3.1). We argue each attribute’s security separately, as BrowserOnly and Monitored are independent.
6.1 BrowserOnly
To access a user’s session, $\sf{\mathcal{A}dv}$ could try to get or set cookies using the Cookies API. However, BrowserOnly cookies are not editable from and never returned to the Cookies API. Thus, $\sf{\mathcal{A}dv}$ cannot use an extension to access BrowserOnly cookies via the Cookie Monster.
$\sf{\mathcal{A}dv}$ can also attempt to access cookies in network headers using the WebRequest and DeclarativeNetRequest APIs. Specifically, $\sf{\mathcal{A}dv}$ can target the Cookie and Set-Cookie HTTP headers. When building network requests, CREAM’s Network Service withholds BrowserOnly cookies from extensions only adding them to the request once all extensions have finished executing. Similarly, when parsing server replies, the Network Service strips any Set-Cookie headers with the BrowserOnly attribute before passing the response to extensions. Therefore, an extension can never access or edit BrowserOnly cookies in network traffic.
$\sf{\mathcal{A}dv}$ could forge a malicious cookie/header in the network traffic to mimic a BrowserOnly cookie. However when attaching BrowserOnly cookies to network requests the Network Service attaches them to the front of the header. This ensures the correct cookies are always parsed first. When parsing server responses, the Network Service overwrites any malicious headers when reattaching BrowserOnly cookies to the response. $\sf{\mathcal{A}dv}$ could also attempt an eviction attack. The browser limits how many cookies a website can set. If the website goes over this limit, the Cookie Monster automatically deletes older cookies to make space for the new ones. Using the Cookies API $\sf{\mathcal{A}dv}$ could generate hundreds of cookies to trick the Cookie Monster into deleting the BrowserOnly cookie. Then $\sf{\mathcal{A}dv}$ could safely attach a fake BrowserOnly cookie to the request headers directly. However, CREAM configures the Cookie Monster to never purge BrowserOnly cookies for space. Hence, $\sf{\mathcal{A}dv}$ cannot forge cookies or headers to edit BrowserOnly cookies and $\sf{\mathcal{A}dv}$ cannot use a malicious extension to steal or modify BrowserOnly cookies.
With a content script, $\sf{\mathcal{A}dv}$ can try to access cookies through the webpage. However, the Cookie Monster also excludes BrowserOnly cookies from JS requests.
Finally, an in-network $\sf{\mathcal{A}dv}$ cannot steal or modify BrowserOnly cookies due to TLS, so long as the cookie is also marked Secure.
6.2 Monitored
Similar to the cases above, $\sf{\mathcal{A}dv}$ may try to get and set cookies using the Cookies API. While Monitored cookies are accessible, the Monitored message and changelog are never returned to the Cookies API. This information is only available through the report generation endpoint internal to the Network Service. Regardless of whether the Monitored message is available, $\sf{\mathcal{A}dv}$ cannot generate a valid report thus invalidating the cookie. $\sf{\mathcal{A}dv}$ does not know the server’s secret key to forge a malicious Monitored message. $\sf{\mathcal{A}dv}$ could try to interleave the Monitored attribute of a different cookie instead, however, the name-value pair in the attribute would not match the stolen cookie. Similarly, $\sf{\mathcal{A}dv}$ cannot change the name-value pair due to the authenticated encryption. Even with the correct Monitored message, $\sf{\mathcal{A}dv}$ would not have the browser’s key and thus cannot generate a valid authenticated Monitored report. $\sf{\mathcal{A}dv}$ also cannot alter the expected key in the Monitored message as they again do not have the server’s secret key.
All changes to a Monitored cookie are appended to the cookie’s changelog and sent to the server as part of the authenticated report. Thus, any changes made by $\sf{\mathcal{A}dv}$ are detectable. To avoid detection, $\sf{\mathcal{A}dv}$ must tamper with the changelog. $\sf{\mathcal{A}dv}$ cannot directly remove entries from the changelog as it is append-only. However, the changelog can be cleared if the cookie’s Monitored message is updated or the cookie is deleted and remade (thus making a new changelog). The first option is impossible as extensions cannot create Monitored cookies. In the second case, $\sf{\mathcal{A}dv}$ could remake the cookie without the Monitored attribute, however, no report would be generated invalidating the cookie.
$\sf{\mathcal{A}dv}$ could also try to steal and tamper with cookies in network traffic using the WebRequest and DeclarativeNetRequest APIs. Monitored cookies and their messages are accessible through the Cookie and Set-Cookie headers. However, Monitored cookies stolen by $\sf{\mathcal{A}dv}$ are unusable. $\sf{\mathcal{A}dv}$ could also edit Monitored cookies through these headers. Since these edits are done directly to the header, they would not appear in the changelog. As such, editing the Set-Cookie header would alter the cookie’s attributes, but these altered settings would appear in the authenticated Monitored report. As a consequence, the server can see that the cookie’s settings do not match the expected configuration and detect the changes. Similarly, if $\sf{\mathcal{A}dv}$ tried to revert the changes before the report generation, the reverted changes would be logged and detectable. $\sf{\mathcal{A}dv}$ could also remove the Monitored attribute itself. This would disable the changelog, but no report would be generated, making the cookie invalid. $\sf{\mathcal{A}dv}$ can also edit cookies in the Cookie header, but this header only contains the cookie’s name and value. Changing either would invalidate the cookie as it would no longer match the expected name-value pair in the Monitored message. $\sf{\mathcal{A}dv}$ could also attempt to tamper with the Monitored message itself. Attempts to remove, replace, or alter the cookie’s Monitored message would invalidate the cookie as it would either generate no report or an invalid report. Finally, $\sf{\mathcal{A}dv}$ could try to tamper with the key and Monitored report also included in network requests. However both the key and report are attached to the request post extensions’ accesses (recall Section 5) and thus are not accessible to extensions.
$\sf{\mathcal{A}dv}$ can use a content script to access cookies through JS. Again, while Monitored cookies can be accessed, their Monitored messages and changelog-s cannot. Similarly, any changes made by $\sf{\mathcal{A}dv}$ are appended to the Monitored cookie’s changelog. JS cannot access the changelog or create Monitored cookies. Thus any edits by malicious JS to Monitored cookies are detectable.
$\sf{\mathcal{A}dv}$ can also intercept, inject, and modify HTTP traffic. As such, $\sf{\mathcal{A}dv}$ may try to steal and modify Monitored cookies over the network. If the Monitored cookie also has the Secure attribute it is always encrypted on the network and thus inaccessible. If the cookie is not Secure, it may be visible. In these cases, $\sf{\mathcal{A}dv}$ can steal and modify Monitored cookies and their Monitored messages in the same way as a malicious extension. However, just like with a malicious extension, Monitored cookies are not usable outside the user’s browser and changes made to network headers invalidate the cookie and are detectable through the Monitored report. Unlike extensions, a network $\sf{\mathcal{A}dv}$ can tamper with the Monitored report as it is included in a custom HTTP header. $\sf{\mathcal{A}dv}$ could strip the report from the request altogether, however, the lack of a report would still invalidate the cookie. $\sf{\mathcal{A}dv}$ could also tamper with the changelog and current settings in the report to hide malicious cookie edits. However, these changes would cause the server-side verification to fail. Similarly, $\sf{\mathcal{A}dv}$ cannot re-authenticate the report after tampering as they do not know the browser’s key for the site. Finally, $\sf{\mathcal{A}dv}$ could try to alter the report, authenticate it with a malicious key, and change the expected key in the Monitored message except $\sf{\mathcal{A}dv}$ does not know the server’s secret key. As such, $\sf{\mathcal{A}dv}$ cannot re-encrypt the Monitored message, verification will fail, and the cookie will be invalid. $\sf{\mathcal{A}dv}$ could also try to replay a previous valid report to hide their actions, however, the report includes a timestamp to greatly reduce this possibility
$\sf{\mathcal{A}dv}$ could try to access/change the key attached to the network request. However, the browser’s key is only attached to HTTPS traffic. As such, the key is always encrypted by TLS on the network and inaccessible to $\sf{\mathcal{A}dv}$.
7 Implementation & Evaluation
We implement CREAM by extending Chromium version 122.0.6168.0 with 535 lines of C++ code across 31 files. For evaluation, we deploy the prototype on a Ubuntu 20.04 machine running at 3.70 GHz with 32GB of RAM. To test CREAM we implemented two Node.js test servers: an HTTPS web application and a malicious endpoint. The test web application allows a user to register an account and log in. When logging in, the server generates a session cookie with pertinent cookie attributes for each test. To allow the server to create cookies with BrowserOnly and Monitored attributes, we added them to a fork of the Node.js cookie library. The malicious server acts as a relay for any messages sent by our malicious extensions discussed in section 3.1. Our prototype and test network can be found on our open-source repository [39] while our fork of Node.js’s cookie library can be found in a separate repository [72]. We reemphasize that these new attributes are optional and CREAM remains compatible with websites that choose not to use them.
7.1 Cookie Sizes & Browser Storage
Each attribute requires storing additional cookie data. The BrowserOnly attribute adds one byte of memory to each cookie. The Monitored attribute is larger, requiring 49 additional bytes. This is due to the Monitored message and changelog being a string and vector, respectively. Both are 24-byte long classes (in a 64-bit architecture) plus there is an additional byte for the default empty character in memory. Monitored also relies on symmetric keys. CREAM uses Chromium’s built-in AES implementation to create the keys. As such, each key is 256 bits long.
The Monitored message and changelog entries also affect the size of Monitored cookies. Both fields depend on server-defined attributes, namely the cookie’s name, value, domain, and path. Prior studies have found the average length of a cookie’s name-value pair to be 37 characters [73] and its domain to be 14 characters long [74]. We use these averages and a default path of “/” for our analysis. In the worst case, all of a cookie’s attributes change at once and thus are included in the resulting changelog entry. CREAM logs changes in an encoded format (i.e. “SameSite: Strict” as “SS:2”) to reduce the entry’s size. As such, this would result in a 108-byte entry and a total memory cost of 132 bytes including the string object itself. However, in most cases, we expect fewer attributes to change at once leading to lower log entry sizes. The Monitored message’s size also varies on the server’s behavior. Every message contains the cookie’s name, value, and the browser’s key at least. In our tests, the key was 32 bytes long resulting in a 131-byte-long encrypted string.
To use either attribute, they must appear in the Set-Cookie header. The BrowserOnly attribute is represented as the 11-byte string “BrowserOnly” whereas Monitored includes both “Monitored=” and the Monitored message. This adds 141 bytes to the header. This is well within Chromium’s 1KB cookie attribute limit [75]. A key and authenticated report are also attached to network requests for Monitored cookies. The browser’s key adds 38 bytes (including the header name) to the request. The attached report contains each Monitored cookie’s sub-report, a timestamp, and an HMAC. A cookie’s sub-report is its current settings, changelog, and Monitored message. The current settings are a 108-byte string. The changelog’s size depends on its entries. Assuming a single worst-case entry (as before), a cookies sub-report is 252 bytes long (with formatting). The timestamp is 15 bytes long and the HMAC is 32 bytes long. Altogether, the report would be 516 bytes long, though this varies greatly on the number of Monitored cookies and their changelog-s. We discuss potential size reductions in Appendix C.
7.2 Operation Run-Times
000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.5BSBOMTime (ms)
(A) Make
002222444466668888BSBOMTime (ms)ExtensionJSNet Service
(B) Get
0022224444666688881010101012121212BSBOMTime (ms)ExtensionJSNet Service
(C) Update
Figure 4: Average time to create, get, and update cookies with CREAM vs. unmodified Chromium.
BO = BrowserOnly; M=Monitored; S=Standard; B=Baseline.
To assess run-time overheads, we measure the time to create, get, and set different cookie types (i.e., “Standard” cookies, Monitored, and BrowserOnly) and compare these times to an unmodified Chromium instance (“Baseline”). This timing depends greatly on the other processes in the browser. As such, we repeat each operation 100,000 times and report the median completion times (Figure 4(C)) to remove the impact of outliers. We measure these times from the Network Service, JS, and extensions. However, creating cookies and all BrowserOnly tests are only measurable from the Network Service. Similarly, all tests are performed synchronously to ensure accuracy.
A cookie’s access time varies on the location of the access. As the Cookie Monster is in the Network Service, accesses from the Network Service are much faster than JS or extensions, which require IPC. The BrowserOnly and Monitored attributes’ additional checks add little overhead to cookie operations. Creating a cookie without either attribute was only 5 $\mu$s slower than the baseline. Fetching a regular cookie from the Network Service, JS, and extensions took 69, 100, and 100 $\mu$s longer than the baseline. Similarly updating a cookie took an additional 100 $\mu$s from both JS and extensions. Curiously, updating a standard cookie through the Network Service was 278 $\mu$s faster than the baseline. Repeating this experiment several times gave us similar results. As CREAM does not optimize cookie operations, this “improvement” reflects the browser’s impact on completion time rather than an actual reduction. Cookie operations in the Network Service are 2$\sim$3 orders of magnitude faster than in JS and extensions (separate processes). As such, operations inside the Network Service are much more impacted by the execution of other processes in the browser likely leading to the discrepancy seen.
BrowserOnly cookies also present a small overhead, requiring 3, 71, and 0 more $\mu$s to be created, retrieved, and updated, respectively. Monitored cookies take an additional 56 $\mu$s to create. Fetching a Monitored cookie through the Network Service, JS, and extensions took an additional 68, 300, and $-100$ $\mu$s. Again the improvement seen when fetching through extensions reflects the impact of other processes and (in this case) IPC on execution time. Updating Monitored cookies took an additional 18, 1,000, and 1,000 $\mu$s respectively. The majority of this overhead stems from updating the cookie’s changelog. Due to this, we also test how varying the numbers of log entries affects a Monitored cookie’s update time. In this test, we timed how long it took to update the cookie 10 times and found the median to be 7,247 $\mu$s. In the original test, updating the cookie once had a median time of 708 $\mu$s. This reflects a linear increase (variability due to background browser operations) indicating that previous log entries do not impact the update time of subsequent entries.
11111.51.51.51.522222.52.52.52.53333BSBOMTime (ms)RequestResponse
(A) Req./Response Handling
000.020.020.020.020.060.060.060.060.10.10.10.1Empty Log1 Entry10 Entries5 CookiesTime (ms)
(B) Report Generation
Figure 5: Latency of (A) different cookie types and (B) Monitored reports for varying number of log entries/cookies.
BO = BrowserOnly; M=Monitored; S=Standard; B=Baseline.
7.3 Server-Side
The BrowserOnly attribute defines how cookies are accessed. Thus BrowserOnly cookies require no additional processing on the server. Monitored cookies, on the other hand, require changes to how servers create and parse cookies. When creating a Monitored cookie, the server must encrypt the Monitored message. Using AES-256-CBC, we found the average time to encrypt to be 6.7 $\mu$s. Monitored cookies also introduce new server-side checks to ensure the report and cookie are valid. While the report verification is always the same, cookie validation depends on the server’s policy. In our experiments, we checked that the changelog was empty and the cookies settings matched an expected configuration. These checks added 3,419 $\mu$s to the server-side processing time.
7.4 Latency
We also measure the time taken to build requests and parse responses with different cookies. For this, we timed how long it took the URLRequestHttpJob to handle the message as only this portion of the network stack changed. Timing this portion alone also helps reduce noise caused by the rest of the browser. Unlike the previous tests, timing the network stack’s behavior required reloading the test page for each sample. As such, we only gathered 10,000 samples to calculate the median in these experiments (depicted in Figure 5(A)).
Parsing responses in CREAM requires BrowserOnly cookies be stripped from responses before any extensions run. Due to this, parsing responses with a BrowserOnly cookie took 133 $\mu$s (4.5%) longer than the baseline. Parsing responses with either a standard or a Monitored cookie has no new behavior. As such, parsing either was on par with the baseline, 14 and 21 $\mu$s faster respectively. Similar to the individual cookie operations, this improvement is due to variability in browser processes. When building the request BroswerOnly cookies must be withheld until after extensions run, Monitored reports must be generated, and the browser key must be added to the headers (if an HTTPS request). When using standard cookies only, CREAM was 55 $\mu$s (3.2%) slower than the baseline despite the only new behavior being attaching the key which took 8 $\mu$s on average. Using BrowserOnly and Monitored cookies took 58 and 220 $\mu$s (3.4% and 12.8%) longer than the baseline. Monitored cookies additional overhead stems from the report generations. As such, we also timed the report generation itself (Shown in Figure 5(B)). Generating a report with no changes took 47 $\mu$s on average. Reports with 1 and 10 changelog entries took 49 and 62 $\mu$s respectively. Finally generating a report with 5 Monitored cookies took 107 $\mu$s on average. This increase comes from running the sub-report generation process multiple times.
8 Privacy Considerations
CREAM’s design aims to strike a good balance between security and privacy. Below we discuss privacy-related considerations.
Private/Incognito Browsing. To limit tracking cookies [76, 77, 78], browsers restrict which contexts cookies can be accessed from. While we discuss the Cookie Monster as a single entity, browsers have different Cookie Monsters for different browsing contexts. For example, Incognito browsing has a separate Cookie Monster from standard browsing. Thus, Incognito cookies are not accessible during standard browsing (and vice-versa). As discussed in Section 2.3, cookies can also be partitioned to restrict their accessibility by embedding context. As CREAM works on individual cookies, it does not inhibit and remains compatible with these privacy features. Further, CREAM-protected cookies remain manageable by users via the normal browser management interfaces.
Privacy-Enhancing Extensions. Some extensions seek to protect user privacy. Among other checks, they often use the Cookies API to search for and remove known tracking cookies from the browser/requests. As the BrowserOnly attribute hides cookies from extensions, it could be leveraged by servers to avoid privacy-enhancing extensions. As a middle ground, it is possible to allow users to selectively disable the BrowserOnly attribute for specific trusted extensions. This opt-out approach would protect the user by default while allowing discretionary disablement. That said, we note that, even without the BrowserOnly attribute, user tracking is still possible through techniques such as browser fingerprinting [79, 80] which are orthogonal to our goal. Further, without approaches to protect cookies, session cookies can be stolen/used, leading to accounts (and all sensitive information therein) being compromised, i.e., a very strong privacy violation. Without browser-based controls such as CREAM, the latter is possible irrespective of privacy-enhancing extensions.
Tracking and Fingerprinting. Monitored cookies rely on cryptographic keys to tie a cookie to a browser. However, as keys are generated per site, they cannot be used to track the user across different domains. It is also possible to generate keys per origin to limit tracking within a single site, e.g., similar to browsers’ origin isolation [81]. We note that, as session cookies contain the user’s authenticated ID, they identify the user. Thus, irrespective of CREAM, while a single session cookie is only valid for a single session, multiple session cookies can still track the user across sessions. However, it is also possible to update/refresh the keys for each session to prevent their use for cross-session tracking. These keys could also allow individual sites to track users across browsing contexts. This can be prevented by generating keys per site and context rather than per site alone. However, this would incur additional runtime and storage costs.
9 Related Work
Given our focus, this section discusses prior work on cookie security. Additional discussion of prior work on extension security is provided in Appendix D.
Cookies are vital to the modern web. As such, their adoption, usage, and vulnerabilities have been well-studied [8, 82, 83, 6, 84]. Their importance has also led to several attempts to mitigate these vulnerabilities. Some prior work introduces additional entities external to the browser to protect session cookies [7, 85, 86]. Often using proxies, these controls store session cookies outside the browser to prevent hijacking [86] or associate them with a companion cookie to mitigate fixation attacks [7]. While effective, external storage only works for cookies that are not used within the browser. Additionally, companion cookies require doubling a website’s cookies. As browsers limit the number of cookies a single site can make [5], companion cookies effectively half the cookies available to a site.
Session Protection is a similar external control offered by Flask-Login [87] (Flask’s session management module). When enabled, the server hashes the user’s IP address and user agent (browser) for each session cookie. These hashes are stored on the server and link the cookies to specific hosts. Then Flask recalculates the hash for all requests with a session cookie and compares it against the saved hash. In the strictest configuration, a hash mismatch invalidates the session. However, by default, a mismatch only marks the session as stale. In this state, the user can still access the site but must re-authenticate for any action that requires a "fresh" session. Further, Flask also supports a separate "Remember Me" token to save and restore sessions. This token is not protected by Session Protection and thus can be stolen and used to bypass Session Protection altogether. Due to this, the Flask-Paranoid extension [88] extends Session Protection to the Remember Me token as well. Regardless, while Session Protection can detect cookie theft, it cannot accurately determine if a cookie was stolen as benign actions (i.e., IP address changes) also violate the control. Thus stolen cookies could still provide access to most of the session.
Other proposals use extensions to bolster the security of cookies [4, 89, 90]. CookieExt [4] detects session cookies in incoming requests and attaches the Secure and HttpOnly attributes to them. Serene [89] maintains a database of session cookie values and ensures the value of outbound cookies are in the database. As discussed in Section 1, malicious extensions can bypass existing cookie attributes and reattach cookies to requests. Thus the aforementioned extension-based approaches cannot protect against other extensions.
Many controls change cookies themselves [91, 2, 92, 93]. Signed cookies contain a cryptographic signature in the cookie’s value to support integrity-checking [92] and access-control [93]. Origin cookies [91] add a new “Origin" attribute that greatly limits a cookie’s accessibility to only the origin that set it. One-Time Cookies (OTC) [2] generate a unique token per request rather than per session. OTCs are then tied to a request with a MAC and verified with a shared session key. This session key and any other information needed for verification are stored in a session ticket, encrypted by the server’s secret key. These features limit how cookies are accessed, but do not protect against and still allow theft by malicious extensions. While OTCs reduce the usability of a stolen cookie to a single request, they are enabled through HTTP headers. Thus, malicious extensions can disable OTCs outright and force standard cookie authentication.
Finally, similar to CREAM, some approaches implement controls directly in the browser [94, 95, 96]. Origin-Bound Certificates (OBCs) [94] use on-demand self-signed certificate generation to allow for mutually authenticated TLS connections. Servers can then associate data such as cookies with an OBC tying it to a single device. The Token Binding standard [95] built upon OBCs, replacing the self-signed certificates with asymmetric keys negotiated during the TLS handshake. The generated Token Binding key and a signature (issued with the corresponding private key) are included in subsequent requests [97]. The Token Binding key is then used as an ID in cookies to tie them to the channel and device. Chrome also recently proposed Device Bound Session Cookies (DBSC) [96]. DBSC proposes using Trusted Platform Module (TPM) keys to tie cookies to a device. In DBSC, the server creates a TPM key on the user’s device. The browser then uses this key to generate a cryptographic proof of identity and sends it to the server. Once verified, the server generates short-lived session cookies to use in future requests. When these cookies expire, DBSC uses the established key to re-authenticate with the server and get fresh cookies. Similar to the Monitored attribute, these controls prevent cookie theft by linking them to a single device. However, none of them address malicious extensions. In Token Binding, both the cookies and signature are included in request headers. As such, both can be stolen by a malicious extension and used to impersonate the user. Similarly, DBSC only changes how session cookies are generated by the server. Thus once the cookies are generated, an extension can still steal them to gain access to the user’s account. DBSC also greatly reduces the lifetime of session cookies, but this only means an adversary needs to steal the cookies more frequently to maintain account access. Also different from CREAM, none of these controls address cookie integrity violations.
10 Conclusion
We proposed CREAM: a modification to the web browser’s trust model to protect cookies from malicious extensions. CREAM introduces two new cookie attributes: BrowserOnly and Monitored. BrowserOnly prevents cookies from being accessed by extensions and JS outright. Monitored maintains accessibility but ties the cookie to a single browser, logs all changes made to the cookie, and sends an authenticated report of changes to the server alongside the cookie. This allows the server to detect and refuse stolen and tampered cookies. We provide an open-source prototype of CREAM built atop Chromium [39] and evaluate it to show that both the BrowserOnly and Monitored attributes incur little overhead.
Appendix A Scraping Browser Extensions
To analyze Chrome’s and Firefox’s extensions we performed a series of iterative scrapes. We began by manually scraping both the Chrome Web Store and AMO, collecting each extension’s name, user count, and URL. Both extension repositories divide extensions into several overlapping categories. Thus, for the AMO, we filtered extensions by category and iterated over all matching results. Doing this, we pulled an initial set of 35,098 Firefox extensions. Scraping the Web Store proved more difficult as search results were loaded dynamically. This required continually pressing a button labeled "load more" to view more results. Further complicating the process, after 50 presses, the webpage would refresh and revert to the initial set of results. Due to this, it was impossible to scrape all Chrome extensions from the search results directly. Instead, we used JS to grab an initial list of 30,473 extension URLs. Then, while scraping the name and user count from each URL, we also pulled any URLs included in the “related extensions” portion of the page. Any previously unseen URLs were added to the scrape. When finished, we found 69,203 Chrome extensions in total.
Compared to prior work this number was low. Between 2018 and 2022, several studies stated the Web Store contained around 130,000 extensions [98, 99, 13] while Chrome themselves reported having over 180,000 [100]. Chrome’s larger estimate likely included both Themes and Chrome Apps. We discuss Themes and Chrome Apps in greater detail later in this section. Due to the discrepancy, we identified a prior scrape performed in 2021 containing 130,445 extension URLs [101] and compared our results against it. From this comparison, we discovered 90,152 previously unseen URLs, however only 20,877 of them still worked presently.
Despite increasing our total Chrome extensions to 90,080, this was still relatively low. Another study on privacy inconsistencies in extensions [98] used sitemaps to discover extensions. Using the sitemap in [102], we found 172,386 extensions. Interestingly, when comparing these results against our prior scrapes, 2,151 previously discovered extensions were missing. This suggests they had been removed from the Web Store between scrapes. We also found AMO’s sitemap [103] and re-scrapped it resulting in 533,768 extensions. Similar to Chrome, 1,170 previously identified extensions were not in the sitemap and were likely removed.
Table 4: Breakdown of manifest versioning, host permissions, and access to cookies
Browser Manifest-Version Host Permissions APIs
V2 V3 $<$all_urls$>$ https://*/* http://*/* Cookies WebRequest WebRequestBlocking
Android 20 2 18 0 0 2 14 13
Chrome Apps 12094 62 455 11 38 15 27 13
A.1 Categorizing extensions
As discussed, we identified 172,386 Chrome and 533,768 Firefox extensions during our scrape. However, the majority of these extensions are "Themes". Themes are purely cosmetic and contain no JS or HTML [104]. For this reason, we exclude Themes from our analysis. Similarly, 16,380 Chrome extensions were unlabeled. We believe these to be Chrome Apps, a now deprecated subclass of extensions that behave similarly to native applications [105]. Though deprecated, Apps are still currently available on Enterprise and Education ChromeOS builds [106]. Hence, we do analyze the Chrome App’s permissions but separately from our main discussion. 22 Firefox extensions were labeled as Android. Given the distinction, we again analyze their permissions separately from the main discussion. Filtering out these categories we were left with 122,951 Chrome and 37,783 Firefox extensions.
A.2 Scraping Manifests
With our dataset of 215,941 extensions (including Chrome Apps and Android variants), the final step was to scrape each extension’s manifest file. Using an online extension viewer [107], we entered each extension’s URL, opened its files, and downloaded its manifest. Despite splitting the dataset across 10 threads, the initial execution took about 36 hours. We then repeated the process for all URLs which resulted in an error with more time between each action. These later runs helped remove false errors due to rate limiting and provided more time for larger extensions to load. When complete, we downloaded 119,152 Chrome and 37,631 Firefox extension manifests. Attempts to scrape the remaining manifests resulted in various 400 HTTP errors, suggesting they were private or removed from the repository.
Appendix B Chrome Apps & Android Analysis
As discussed in Section 3.2, we discovered 16,380 Chrome Apps and 22 Firefox for Android extensions. From this, we pulled 12,156 App and 22 Android manifests and analyzed them for their version, host permissions, and to see if they have access to cookies. A summary of our findings can be found in Table 4.
Nearly all manifests analyzed (99.5%) were written using manifest V2. This is likely due to a lack of incentive to change formats as Firefox currently has no plans to deprecate the format and Chrome Apps will be fully deprecated in 2025 [105]. Looking at host permissions, only 4.3% of manifests declared broad host permissions. However, most Android extensions (81.8%) had “$<$all_urls$>$" permissions. Only 4.1% of Chrome Apps had any broad host permissions. This lack of permission may be due to Chrome Apps performing more like native applications thus running more locally rather than needing to access many websites.
We saw similar low numbers concerning access to cookies. Only 17 manifests (0.1%) had the Cookies permission. Similarly, only 41 (0.3%) used the WebRequest API with 26 (0.2%) also having access to the WebRequestBlocking permission. While low, 63.6% of Android extensions had the WebRequest permission, and all but one of these also declared the WebRequesBlocking permission. Despite this, manifests with the Cookies, WebRequest, and WebRequestBlocking API affected a combined 2,245,591, 14,699,240, and 13,622,263 users respectively. No manifest declared the DeclarativeNetRequest permission. We further analyzed which manifests that could access cookies also had “$<$all_urls$>$" host permissions. Table 5 summarizes these findings. Both Android extensions with the Cookies permission could access all cookies in the browser. This was less common in Chrome Apps with only 20% of those with Cookies permissions able to access all cookies. 92.9% of Android extensions and 48.1% of Chrome Apps with WebRequest permissions can see all traffic in the browser. Similarly, all of these Android extensions and 60% of these Chrome Apps also declare the WebRequestBlocking permission.
Table 5: Manifests with ”$<$all_urls$>$” and cookie access
Browser Cookies WebRequest WebRequestBlocking
Android 2 13 13
Chrome Apps 3 13 8
Appendix C Additional Discussion & Limitations
Reduced Memory Overhead: As mentioned in Section 7.1, both the Monitored changelog and report can grow rapidly for cookies with frequent changes. Thus one direction for future work is to minimize the size of these values. There are several ways this reduction may be achieved. In our design, both the Monitored log entries and message are C++ string objects. These objects add 24 bytes of overhead to each cookie and each changelog entry. Thus switching to another representation such as a basic character pointer could help reduce some of the storage overhead. Similarly, we already encode the changelog entries and current settings to minimize their size. However, as they are both strings it may be possible to use other string compression algorithms to reduce their size further. Finally, it may be possible to reduce which attributes are logged. Some attributes such as HttpOnly prevent cookies from being stolen, but a stolen Monitored cookie cannot be used outside the user’s browser. As such it may not be necessary to log changes made to these attributes.
Several possible server-side changes could also reduce the memory overhead of Monitored cookies. Each Monitored cookie’s message contains an encrypted copy of the browser’s key for the domain. However all Monitored cookies for the same domain share the same key. Thus it may be possible to store the encrypted key in a subset of cookies that cover all possible requests. This would reduce the Monitored message and report size of many Monitored cookies as they would no longer need to store a duplicate encrypted key. Further, if the server stored the browser’s key itself (similar to proposals like DBSC [96] or frameworks like Flask [87]) there would be no need for the Monitored message at all. However, this would trade browser overhead for server overhead. Finally, it may be possible to create a new server endpoint (also similar to DBSC) to receive Monitored reports. This would move the reports from a header and into the request body which has fewer size limitations. However, all of these changes would require increased server-side support.
Further Extension Analysis: During our analysis of extensions, we specifically looked for which extensions could access cookies. However, during the process, we noted some other interesting results. During the initial scrape, we identified several nearly identical extensions. These extensions had identical names, highly similar descriptions and assets, and the majority of their code was the same. However, it is unclear why some many duplicates exist and how they differ. Similarly, while Chrome and Firefox both support the WebExtension APIs they do not always share the same implementation. Despite this many extensions work across browsers. Due to these observations, another interesting direction for future work is further analysis of these extensions to determine how they differ from each other, across browsers, and what effect this may have on the security of the browser.
Appendix D Extension Security: Extended Related Work
Given their popularity and access to powerful APIs, several studies have analyzed extensions in terms of the APIs they commonly use, how they impact other controls, and how users understand their impacts [99, 98, 13, 30]. Additionally, prior work also investigates the privacy threats of extensions, such as tracking user behavior and interests [108]. As such many works investigate extension fingerprinting techniques [109, 110] and ways to prevent it [111, 112].
As mentioned in Section 1, extensions can be exploitable yielding extension permissions to $\sf{\mathcal{A}dv}$. Many static analysis techniques and browser modifications have been proposed to help detect/mitigate extension vulnerabilities [34, 113, 114, 33]. Several studies also acknowledge the threat posed by malicious extensions focusing on detection [25, 115, 24, 116, 117, 118, 26] whereas fewer prior attempts aim to prevent malicious extension actions. WebEnclave [35] is an extension that creates an isolated enclave on a website to hide secure elements such as login prompts from other extensions. Similarly, WRIT [36] uses service workers to attach tokens to JS events (i.e. mouse clicks) to verify if they were generated by the user rather than an extension. Picazo-Sanchez et. al. [37] create a browser extension to monitor changes made by each extension, remove them before the next extension runs, and re-add all the changes after all extensions have executed. The latter prevents a malicious extension from abusing execution orders to learn about/tamper with other extensions. These works focus on website interactions rather than broader browser APIs like Cookies and WebRequest.